OpenStudy (anonymous):

What is the fifth term of a sequence whose first term is 5 and whose common ratio is 3? I'm confused

OpenStudy (anonymous):

\[5,5\times 3,5\times 3^2,5\times 3^3,...\]

OpenStudy (kamille):

where is a formula \[a _{n}=a _{1}*r ^{n-1}\] you have \[a _{5}=5*3^{5-1}\] all you need to do is calculate now;)

OpenStudy (zehanz):

First term 5. Second term 5*3 Third term 5*3^3 Fourth term 5*3^3 Fifth term 5*3^4 = 5*81 = ...

OpenStudy (anonymous):

405

OpenStudy (zehanz):

Right!